### v1

@@ 3.3 CoSE Module
In the original UNet, each decoder stage consists of only two groups of (CNN, batch normalization (BN), and rectified linear unit (ReLU) activation), as in Figure 1(e). The inference rules are illustrated below:
@@
For medical images, the target regions are usually intersected with complicated background organs, presenting similar textures and shapes. It is required to improve the representation of critical features and suppress the impact of irrelevant background features. To this end, in this work, the channel attention mechanism (SENet) is introduced to the original decoder stages to capture the interdependency among channels by integrating the information of channel context. The SENet can adaptively adjust the importance of each feature channel through learnable attention weights, thus highlighting the important regions and suppressing the background regions.
As shown in Figure 1(b), the CoSE module is constructed by a convolution block and a SENet, in which corresponding residual connections are added before the convolution block and the SENet, respectively. The residual connection directly applies the addition operation to transmit the learned features into deeper layers, which can effectively solve the gradient vanishing and explosion by providing additional paths to enhance the information propagation. The mentioned process can be expressed as the equations:
@@
where SENet is implemented by referring to (Hu et al. (2018)).
@@ 3.4 MLF Module
In the proposed model, the pre-trained encoder is applied to extract the features from input images by reducing the feature resolution with higher channels, while the decoder generates the pixel-wise segmentation masks by recovering the resolution to the raw image. In the original UNet, the interpolation algorithm is applied to implement the up-sampling operation, limiting the feature interactions among different scales and losing key details in the segmentation results.
@@
To address this issue, in this work, an MLF module is proposed to fuse the learned features in decoder stages with different scales, as in Figure 3. The primary purpose is to enhance feature details by interaction and fusion operations, which helps the model to better identify different regions with similar features. The output feature maps from decoder stages 1-4 serve as the MLF inputs, which are up-sampled by 2, 4, 8, and 16 to generate image size-liked masks. Then, the channel number of the up-sampled feature maps is mapped to the class number relevant to the task by a point-wise convolution. The inference rules are listed as:
@@ 3.5 Loss Function
To enhance the model convergence, the multi-stage feature mixing loss aggregation (MUTATION) method proposed by MERIT (Rahman and Marculescu (2024)) is introduced in this work. To be specific, for the feature maps ($Y_{1}$, $Y_{2}$, $Y_{3}$, $Y_{4}$) from the MLF module, a total of 15 ($12^{4}-$ 1) nonempty subsets are first obtained, based on which 15 predicted masks are generated by element-wise addition on the feature maps in each set. Consequently, for each mask in the 15 predictions, the $L$ in Equation 20 is calculated over the ground truth to measure multi-stage prediction errors. In addition, the loss between the MLF output and the ground truth is also considered to formulate the final loss, as in:
@@ 4.1 Datasets and evaluation metrics
Synapse multi-organ segmentation dataset: 30 abdominal CT scans are included in the Synapse (Landman et al. (2015)), with a total of 3779 axial contrast-enhanced abdominal CT images, which is provided by the MICCAI 2015 Multi-Atlas Abdomen Labeling Challenge. Each CT scan consists of 85-198 slices of 512 × 512 pixels, and the voxel spatial resolution is ([0:54-0:54] × [0:98-0:98] × [2:5-5:0])$m{m^{3}}$. Following TransUNet (Chen et al. (2021)), the dataset is randomly divided into 18 scans (2212 axial slices) for training, and 12 for validation. A total of 8 anatomical structures are segmented, including the aorta, gallbladder, left kidney, right kidney, liver, pancreas, spleen, and stomach. DICE scores and 95% Hausdorff Distance (95HD) are used as the evaluation metrics in the experiments on this dataset.
Table 1: Comparative analysis of model performance on Synapse dataset for organ segmentation. Organ abbreviations: GB (gallbladder), KL (left kidney), KR (right kidney), PC (pancreas), SP (spleen), SM (stomach). Only Dice scores are reported for individual organs. High DICE scores and low HD95 scores mean better performance. The best result is highlighted in bold, and the second-best is highlighted with an underline. (TransCASCADE: 123.47M; Cascaded MERIT: 147.86M; Small FIF-UNet: 86.91M; Tiny FIF-UNet: 38.31M)
Methods Average Aorta GB KL KR Liver PC SP SM
DICE↑ HD95↓
CNN UNet 70.11 44.69 84.00 56.70 72.41 62.64 86.98 48.73 81.48 67.96
AttnUNet 71.70 34.47 82.61 61.94 76.07 70.42 87.54 46.70 80.67 67.66
ViT SwinUNet 79.13 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
TransDeepLab 80.16 21.25 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
MISSFormer 81.96 18.20 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
Hybrid TransUNet 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
SSFormerPVT 78.01 25.68 82.78 63.74 80.72 78.11 93.53 61.53 87.07 76.61
PolypPVT 78.08 25.61 82.34 66.14 81.21 73.78 94.37 59.34 88.05 79.40
MT-UNet 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
HiFormer 80.29 18.85 85.63 73.29 82.39 64.84 94.22 60.84 91.03 78.07
PVT-CASCADE 81.06 20.23 83.01 70.59 82.23 80.37 94.08 64.43 90.10 83.69
CASTformer 82.55 22.73 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
TransCASCADE 82.68 17.34 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.52
Cascaded MERIT 84.90 13.22 87.71 74.40 87.79 84.85 95.26 71.81 92.01 85.38
Ours Tiny FIF-UNet 84.40 19.89 89.19 74.56 85.22 81.39 95.19 70.69 92.11 86.42
Small FIF-UNet 86.05 15.82 89.49 76.15 88.23 86.26 95.87 74.14 91.31 86.97
Automated cardiac diagnosis challenge: The ACDC dataset (Bernard et al. (2018)) consists of 100 cardiac MRI scans collected from different patients, provided by the MICCAI ACDC challenge 2017. Each scan contains three organs: the right ventricle, left ventricle, and myocardium. Following TransUNet (Chen et al. (2021)), 70 cases (1930 axial slices) are used for training, 10 for validation, and 20 for testing. DICE score serves as the evaluation metric in the experiments on this dataset.
@@ 4.2 Implementation details
The pre-trained MaxViT from (Tu et al. (2022)) serves as the encoder of the proposed model, with the input resolution of 256×256 and attention window size of 7×7. To consider the final performance, both the small and tiny MaxViT architectures are applied to conduct the experiments, i.e., Small FIF-UNet and Tiny FIF-UNet. To enhance the diversity of the training samples, random rotation, and flipping strategies are performed to augment the raw data (Chen et al. (2021)). The model is trained using AdamW (Loshchilov and Hutter (2017)) optimizer with the learning rate of 1e-4 for 400 epochs, applying the weight decay of 1e-4. The batch size of 16 is used for Synapse and ACDC. Following (Tu et al. (2022)), the loss weights ${\lambda_{1}}$ and ${\lambda_{2}}$ are set to 0.7 and 0.3, respectively. The proposed model is implemented using Pytorch 2.2.2 and all experiments are conducted on a single NVIDIA TITAN RTX GPU with 24GB of memory.
4.3 Results and Discussions
@@ 4.3.1 Experimental results on Synapse dataset
The experimental results on the Synapse multi-organ dataset are reported in Table 1, including the proposed model and other selective baselines. A total of 14 comparative baselines are selected to evaluate the model performance, as in three categories, CNN-based models (i.e., UNet (Ronneberger et al. (2015)), AttnUNet (Oktay et al. (2018))), ViT-based models (i.e., SwinUNet (Cao et al. (2022)), TransDeepLab (Azad et al. (2022a)), MISSFormer (Huang et al. (2021))) and hybrid CNN-Transformer models (i.e., TransUNet (Chen et al. (2021)), SSFormerPVT (Wang et al. (2022)), PolypPVT (Dong et al. (2021)), MT-UNet (Wang et al. (2022)), HiFormer (Heidari et al. (2023)), PVT-CASCADE (Rahman and Marculescu (2023)), CASTformer (You et al. (2022)), TransCASCADE (Rahman and Marculescu (2023)), Cascaded MERIT (Rahman and Marculescu (2024))).
@@
As shown in Table 1, the average DICE and HD95 are reported to compare the model performance, as well as the DICE score for the certain 8 classes. In general, the Small FIF-UNet achieves the highest average DICE of 86.05% (primary metric), which significantly outperforms all the selective baselines (1.15% absolute improvement over the best baseline). Specifically, the Small FIF-UNet has the ability to harvest the best performance for 7/8 classes, confirming the performance superiority over baselines. Meanwhile, the proposed model also obtains the second-best HD95 measurement.
Compared to classical ViT-based SwinUNet and hybrid TransUNet, the Small FIF-UNet improves the performance in terms of average DICE by 6.92% and 8.57%, respectively. More importantly, the proposed FIF-UNet achieves significant performance improvements in both small and hard-to-segment organs. Specifically, the proposed model improves by 0.44% and 1.41% for small organs (KL and KR) over the best baseline, respectively, while improving by 1.75% and 2.33% for hard-to-segment organs (GB and PC).
In addition, we also consider the model complexity. Compared to the two competitive baselines TransCASCADE and Cascaded MERIT, the Small FIF-UNet can achieve the best performance with only 86.91M parameters (36.56M and 60.95M lower than baselines). It is worth noting that the proposed Tiny FIF-UNet achieves the third-best performance of 84.4% average DICE among baselines, comparable to the second-best Cascaded MERIT, with only 38.31M trainable parameters (109.55M lower).
@@ 4.3.2 Experimental results on ACDC dataset
Table 2: Comparative analysis of model performance on ACDC dataset for organ segmentation. Organ abbreviations: RV (right ventricle), Myo (myocardium), LV (left ventricle). Only Dice scores are reported for individual organs. The best result is highlighted in bold, and the second-best is highlighted with an underline. (Parallel MERIT: 147.86M; Small FIF-UNet: 86.91M; Tiny FIF-UNet: 38.31M)
Methods Avg DICE↑ RV Myo LV
CNN R50 AttnUNet 86.75 87.58 79.20 93.47
R50 UNet 87.55 87.10 80.63 94.92
ViT SwinUNet 90.00 88.55 85.62 95.83
MISSFormer 90.86 89.55 88.04 94.99
Hybrid TransUNet 89.71 88.86 84.53 95.73
MT-UNet 90.43 86.64 89.04 95.62
PVT-CASCADE 91.46 88.90 89.97 95.50
TransCASCADE 91.63 89.14 90.25 95.50
Cascaded MERIT 91.85 90.23 89.53 95.80
Parallel MERIT 92.32 90.87 90.00 96.08
Ours Tiny FIF-UNet 92.37 91.04 90.10 95.96
Small FIF-UNet 92.58 91.30 90.24 96.19
@@
The experimental results of comparative methods on the ACDC dataset are reported in Table 2, in terms of the average DICE score. Similarly, three categories of models are selected as the comparative baselines, including CNN-based models(i.e., R50 UNet (Chen et al. (2021)), R50 AttnUNet (Chen et al. (2021))), ViT-based models (i.e., SwinUNet (Cao et al. (2022)), MISSFormer (Huang et al. (2021))) and hybrid CNN-Transformer models (i.e., TransUNet (Chen et al. (2021)), MT-UNet (Wang et al. (2022)), PVT-CASCADE (Rahman and Marculescu (2023)), TransCASCADE (Rahman and Marculescu (2023)), Parallel MERIT (Rahman and Marculescu (2024)), Cascaded MERIT (Rahman and Marculescu (2024))).
As shown in Table 2, the proposed Small FIF-UNet outperforms recent state-of-the-art (SOTA) methods with an average DICE score of 92.58%, over the best baseline Parallel MERIT (92.32%). Specifically, the Small FIF-UNet achieves the best DICE score in RV (91.30%) and LV (96.19%) segmentation and the second best DICE score in Myo (90.24%, only 0.01% inferior) segmentation. In addition, the proposed Tiny FIF-UNet achieves the second-best results (92.37%), while the number of parameters is 109.55M lower than the previous best result.
Notably, the proposed FIF-UNet employs the same pre-trained encoder with Parallel MERIT, our model has the ability to achieve higher performance for all three organs. This can be attributed that the proposed technical modules on both the decoder and skip connections enhance the model performance. In summary, the results not only confirm the effectiveness and efficiency of the proposed model but also support our motivations to improve the skip connections and decoder.

### v2

@@ 3.3 CoSE Module
In the original UNet, each decoder stage consists of only two groups of (CNN, batch normalization (BN), and rectified linear unit (ReLU) activation). The inference rules are illustrated below:
@@
In medical images, target regions often share similar textures and shapes with complex background organs, further providing challenges to distinguish critical features. To address this, the channel attention SENet (Hu et al. (2018)) is integrated into the decoder stages to model channel interdependencies. SENet adaptively recalibrates feature channels via learnable attention weights, enabling the network to extract informative channels related to the target regions while suppressing irrelevant background channels.
As shown in Figure 1(a), the CoSE module is constructed by a convolution block and a SENet, in which corresponding residual connections are added before the convolution block and the SENet, respectively. The residual connection directly applies the addition operation to transmit the learned features into deeper layers, which can effectively solve the gradient vanishing and explosion by providing additional paths to enhance the information propagation. The mentioned process can be expressed as the equations:
@@
This design is intuitively motivated by enhancing channel-wise feature selection within the decoder to improve the representation of discriminative features, which is critical for highlighting target regions from complex backgrounds. The residual connections further ensure stable training and effective feature propagation.
@@ 3.4 MLF Module
In the proposed model, the pre-trained encoder extracts features from input images by progressively reducing spatial resolution while increasing channel depth. The decoder then restores the spatial resolution to generate pixel-wise segmentation masks. In the original UNet, up-sampling is typically performed using simple interpolation, which limits effective interactions between features of different scales and often results in loss of fine-grained details important for accurate segmentation.
@@
To address this issue, a MLF module, as illustrated in Figure 4, is proposed to explicitly fuse features from different decoder stages at multiple scales. The MLF module enables richer feature interactions across scales through learnable fusion operations to preserve spatial details. The goal of the MLF design is that integrating multi-scale information allows the network to combine coarse semantic context with fine spatial details, resulting in more precise and robust segmentation outputs.
The output feature maps from decoder stages 1-4 serve as the MLF inputs, which are up-sampled by 2, 4, 8, and 16 to generate image size-liked masks. Then, the channel number of the up-sampled feature maps is mapped to the class number relevant to the task by a point-wise convolution. The inference rules are listed as:
@@ 3.5 Loss Function
To enhance the model convergence, the multi-stage feature mixing loss aggregation (MUTATION) method proposed by MERIT (Rahman and Marculescu (2024)) is introduced in this work. To be specific, for the feature maps ($Y_{1}$, $Y_{2}$, $Y_{3}$, $Y_{4}$) from the MLF module, a total of 15 ($2^{4}-1$) nonempty subsets are first obtained, based on which 15 predicted masks are generated by element-wise addition on the feature maps in each set. Consequently, for each mask in the 15 predictions, the $L$ in Equation 6 is calculated over the ground truth to measure multi-stage prediction errors. In addition, the loss between the MLF output and the ground truth is also considered to formulate the final loss, as in:
@@ 4.1 Datasets and evaluation metrics
[…]
Table 2: Comparative analysis of model performance on Synapse dataset for organ segmentation. Organ abbreviations: GB (gallbladder), KL (left kidney), KR (right kidney), PC (pancreas), SP (spleen), SM (stomach). Only Dice scores are reported for individual organs. High Dice scores and low HD95 scores mean better performance. The best result is highlighted in bold, and the second-best is highlighted with an underline.
Methods Average Aorta GB KL KR Liver PC SP SM Params FLOPs
DICE↑ HD95↓ (M) (G)
CNN UNet 70.11 44.69 84.00 56.70 72.41 62.64 86.98 48.73 81.48 67.96 31.04 54.8
AttnUNet 71.70 34.47 82.61 61.94 76.07 70.42 87.54 46.70 80.67 67.66 34.88 66.64
nnU-Net 82.63 - 91.28 64.06 83.57 81.31 94.53 73.67 88.60 84.05 31.06 78.38
ViT SwinUNet 79.13 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60 27.17 5.92
TransDeepLab 80.16 21.25 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40 21.14 15.75
MISSFormer 81.96 18.20 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81 42.46 7.28
Hybrid TransUNet 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62 105.28 25.41
SSFormerPVT 78.01 25.68 82.78 63.74 80.72 78.11 93.53 61.53 87.07 76.61 - -
PolypPVT 78.08 25.61 82.34 66.14 81.21 73.78 94.37 59.34 88.05 79.40 25.11 10.10
MT-UNet 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81 79.08 44.87
HiFormer 80.29 18.85 85.63 73.29 82.39 64.84 94.22 60.84 91.03 78.07 25.51 13.33
PVT-CASCADE 81.06 20.23 83.01 70.59 82.23 80.37 94.08 64.43 90.10 83.69 35.28 7.62
CASTformer 82.55 22.73 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55 - -
TransCASCADE 82.68 17.34 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.52 123.49 41.03
Cascaded MERIT 84.90 13.22 87.71 74.40 87.79 84.85 95.26 71.81 92.01 85.38 147.86 33.44
Tiny FIF-UNet(ours) 84.40 19.89 89.19 74.56 85.22 81.39 95.19 70.69 92.11 86.42 38.31 11.21
Small FIF-UNet(ours) 86.05 15.82 89.49 76.15 88.23 86.26 95.87 74.14 91.31 86.97 86.91 23.26
Automated cardiac diagnosis challenge: The ACDC dataset (Bernard et al. (2018)) consists of 100 cardiac MRI scans collected from different patients, provided by the MICCAI ACDC challenge 2017. Each scan contains three organs: the right ventricle, left ventricle, and myocardium. Following the official data split in the ACDC challenge and consistent with previous works, the dataset is divided into 70 cases (1930 axial slices) for training, 10 for validation, and 20 for testing. The results reported in this paper are based on the test set to ensure fair comparison and reproducibility. The Dice score is used consistently as the evaluation metric during all stages of training, validation, and testing on this dataset.
@@ 4.2 Implementation details
The pre-trained MaxViT from (Tu et al. (2022)) serves as the encoder of the proposed model, with the input resolution of 256×256 and attention window size of 7×7. To consider the final performance, both the small and tiny MaxViT architectures are applied to conduct the experiments, i.e., Small FIF-UNet and Tiny FIF-UNet. To enhance the diversity of the training samples, we implement three data augmentation strategies: (1) discrete random rotation (90°, 180°, or 270°), (2) random flipping (horizontal or vertical), and (3) continuous random rotation (with angles varying within ±20°). The model is trained using AdamW (Loshchilov and Hutter (2017)) optimizer with the learning rate of 1e-4 for 400 epochs, applying the weight decay of 1e-4. The batch size of 16 is used for Synapse and ACDC. Following (Tu et al. (2022)), the loss weights ${\lambda_{1}}$ and ${\lambda_{2}}$ are set to 0.7 and 0.3, respectively. The proposed model is implemented using Pytorch 2.2.2 and all experiments are conducted on a single NVIDIA TITAN RTX GPU with 24GB of memory.
4.3 Quantitative Results
This section provides quantitative comparisons of the proposed model against state-of-the-art methods in terms of both segmentation accuracy and computational efficiency. Accuracy is evaluated on the primary Synapse dataset and the ACDC dataset. Computational efficiency is measured and compared using model parameter count (Params) and floating-point operations (FLOPs). The results validate that our model achieves superior accuracy, robust generalization, and a better trade-off between performance and resource consumption. The detailed quantitative results are presented below.
@@ 4.3.1 Experimental results on Synapse dataset
The experimental results on the Synapse multi-organ dataset are reported in Table 2, including the proposed model and other selective baselines. A total of 14 comparative baselines are selected to evaluate the model performance, as in three categories, CNN-based models (i.e., UNet (Ronneberger et al. (2015)), AttnUNet (Oktay et al. (2018)), nnU-Net(Isensee et al. (2021))), ViT-based models (i.e., SwinUNet (Cao et al. (2022)), TransDeepLab (Azad et al. (2022a)), MISSFormer (Huang et al. (2021))) and hybrid CNN-Transformer models (i.e., TransUNet (Chen et al. (2021)), SSFormerPVT (Wang et al. (2022)), PolypPVT (Dong et al. (2021)), MT-UNet (Wang et al. (2022)), HiFormer (Heidari et al. (2023)), PVT-CASCADE (Rahman and Marculescu (2023)), CASTformer (You et al. (2022)), TransCASCADE (Rahman and Marculescu (2023)), Cascaded MERIT (Rahman and Marculescu (2024))).
@@
As shown in Table 2, the average DICE and HD95 are reported to compare the model performance, as well as the Dice score for the certain 8 classes. In general, the Small FIF-UNet achieves the highest average DICE of 86.05% (primary metric), which significantly outperforms all the selective baselines (1.15% absolute improvement over the best baseline). Specifically, the Small FIF-UNet has the ability to harvest the best performance for 7/8 classes, confirming the performance superiority over baselines. Meanwhile, the proposed model also obtains the second-best HD95 measurement. Compared to classical ViT-based SwinUNet and hybrid TransUNet, the Small FIF-UNet improves the performance in terms of average DICE by 6.92% and 8.57%, respectively.
More importantly, the proposed FIF-UNet demonstrates significant performance improvements, particularly in small and hard-to-segment organs. Compared to the Cascaded MERIT(use the same pre-trained encoder), on two small and relatively underrepresented organs, the Small FIF-UNet achieves improvements of 0.44% on the left kidney (KL) and 1.41% on the right kidney (KR), which can be attributed that the MLF module can enhance feature detail through multi-scale fusion to improve the model’s ability to detect small structures.
For challenging organs with low contrast and ambiguous boundaries, such as the gallbladder (GB) and pancreas (PC), the proposed model achieves even larger performance gains of 1.75% and 2.33%, respectively. These improvements benefit from the combined effects of the CSI module to strengthen semantic interactions between encoder and decoder layers by preserving more informative details, and the CoSE module to enhance the model’s capacity to represent discriminative features. By combining the two blocks, the model provides better anatomical consistency across organs of varying size and difficulty.
In addition, the proposed Tiny FIF-UNet and Small FIF-UNet achieve an excellent balance between segmentation accuracy and computational efficiency. From the efficiency perspective, the Tiny FIF-UNet achieves an average Dice score of 84.40% with only 38.31M parameters and 11.21 GFLOPs, outperforming models with comparable or even higher complexity (i.e., AttnUNet (71.70%, 34.88M, 66.64G) and HiFormer (80.29%, 25.51M, 13.33G)), demonstrating its superior accuracy-efficiency performance. From the accuracy perspective, the Small FIF-UNet achieves the highest average Dice score of 86.05%, outperforming all baselines including TransCASCADE (82.68%, 123.49M, 41.03G) and Cascaded MERIT (84.90%, 147.86M, 33.44G), while using significantly fewer parameters and FLOPs. The results show that the proposed model can provide better performance at a lower computational cost.
@@ 4.3.2 Experimental results on ACDC dataset
Table 3: Comparative analysis of model performance on ACDC dataset for organ segmentation. Organ abbreviations: RV (right ventricle), Myo (myocardium), LV (left ventricle). Only Dice scores are reported for individual organs. The best result is highlighted in bold, and the second-best is highlighted with an underline.
Methods Avg DICE↑ RV Myo LV Params(M) FLOPs(G)
CNN R50 AttnUNet 86.75 87.58 79.20 93.47 - -
R50 UNet 87.55 87.10 80.63 94.92 - -
nnU-Net 91.59 90.53 89.91 94.33 31.06 78.38
ViT SwinUNet 90.00 88.55 85.62 95.83 27.17 5.92
MISSFormer 90.86 89.55 88.04 94.99 42.46 7.28
Hybrid TransUNet 89.71 88.86 84.53 95.73 105.28 25.41
MT-UNet 90.43 86.64 89.04 95.62 79.08 44.87
PVT-CASCADE 91.46 88.90 89.97 95.50 35.28 7.62
TransCASCADE 91.63 89.14 90.25 95.50 123.47 41.03
Cascaded MERIT 91.85 90.23 89.53 95.80 147.86 33.44
Parallel MERIT 92.32 90.87 90.00 96.08 147.86 33.44
Tiny FIF-UNet(ours) 92.37 91.04 90.10 95.96 38.31 11.21
Small FIF-UNet(ours) $\mathbf{92.58\pm 0.12}$ 91.30 90.24 96.19 86.91 23.26
@@
The experimental results of comparative methods on the ACDC dataset are reported in Table 3, in terms of the average Dice score. Similarly, three categories of models are selected as the comparative baselines, including CNN-based models (i.e., R50 UNet (Chen et al. (2021)), R50 AttnUNet (Chen et al. (2021)), nnU-Net(Isensee et al. (2021))), ViT-based models (i.e., SwinUNet (Cao et al. (2022)), MISSFormer (Huang et al. (2021))) and hybrid CNN-Transformer models (i.e., TransUNet (Chen et al. (2021)), MT-UNet (Wang et al. (2022)), PVT-CASCADE (Rahman and Marculescu (2023)), TransCASCADE (Rahman and Marculescu (2023)), Cascaded MERIT (Rahman and Marculescu (2024)), Parallel MERIT (Rahman and Marculescu (2024))).
As shown in Table 3, the proposed Small FIF-UNet achieves an average Dice score of 92.58±0.12% on the ACDC dataset across eight independent runs, with a confidence range of 0.12. Compared to the best baseline Parallel MERIT (92.32%) with the same pre-trained encoder, the proposed model also achieves 0.26% improvement, and the tight confidence range clearly indicates the greater stability and reliability of our model.
Specifically, the Small FIF-UNet achieves the highest Dice scores on RV (91.30%) and LV (96.19%), and comparable performance for Myo (90.24%), only 0.01% below the best method. These improvements are especially valuable given the challenges of segmenting the right ventricle and myocardium, which often exhibit blurry boundaries and low contrast with the background.
It is believed that the consistent performance improvement can be attributed to the proposed architectural components: the CSI module enhances multi-scale semantic fusion via skip connections; CoSE improves the focus on discriminative regions; and MLF reinforces fine-grained structural details. By integrating them into the proposed model, both feature representation and boundary localization can be enhanced and results across Synapse and ACDC datasets also demonstrate the robust generalization capability across different data modalities.
Similarly, the proposed model achieves the balance between segmentation accuracy and computational efficiency to the Synapse dataset, as that in ACDC. Notably, the proposed Tiny FIF-UNet achieves the second-best average Dice score of 92.37%, closely matching Parallel MERIT (92.32%) while using only 38.31M trainable parameters v.s. 109.55M in Parallel MERIT. Considering that the two models adopt the same pre-trained encoder, this result further validates the strong capability of the proposed modules to enhance feature representation and segmentation performance with significantly lower computational cost.